An aqueous solution containing 6.16 g of lead(II) nitrate is added to an aqueous solution containing 6.62 g of potassium chloride.
what is the balanced chemical equation for this reaction?
what is the limiting reactant?The percent yield for the reaction is 79.7% . How many grams of precipitate is recovered?
How many grams of the excess reactant remain?
1
Expert's answer
2019-10-02T08:39:48-0400

Solution.

The balanced chemical equation for this reaction:

"Pb(NO3)2 + 2KCl = PbCl2 + 2KNO3"

Now we find the limiting reagent through the amount of substance:

"n(Pb(NO3)2) = \frac{m(Pb(NO3)2)}{M(Pb(NO3)2)}"

"n(KCl) = \frac{m(KCl)}{M(KCl)}"

n(Pb(NO3)2) = 0.02 mole

n(KCl) = 0.09 mole

So, Pb(NO3)2 is limiting reagent.

Now we should find m(PbCl2). Since n(Pb(NO3)2) is equal to n(PbCl2), so we find m(PbCl2).

"m(PbCl2) = n(PbCl2) \times M(PbCl2)"

m(PbCl2, theor.) = 5.56 g

The yield of the reaction product is the ratio of the mass of the substance obtained in the experiment (practical) to the mass of the precipitate obtained in theory (theoretical).

"\eta = \frac{m(PbCl2, pract.)}{m(PbCl2, theor.)} \times 100 \%"

To find the mass of sediment in the experiment, we express the practical mass from the formula.

"m(PbCl2, pract.) = \frac{\eta \times m(PbCl2, theor.)}{100 \%}"

m(PbCl2, pract.) = 4.43 g

Answer:

The balanced chemical equation for this reaction:

"Pb(NO3)2 + 2KCl = PbCl2 + 2KNO3"

Pb(NO3)2 is limiting reagent.

m(PbCl2, pract.) = 4.43 g